OpenStudy (anonymous):

on two investments totaling $11,500, Jessica lost 3% on one and earned 7% on the other. If her net annual receipts were $196, how was each investment? ...use the variables x and y to set up a system of equations

OpenStudy (anonymous):

@Hero ?

OpenStudy (anonymous):

@hartnn ?

OpenStudy (anonymous):

@jim_thompson5910 ?

OpenStudy (anonymous):

@satellite73 I really need help with this..anyone?

OpenStudy (anonymous):

we can do this i am sure

OpenStudy (anonymous):

i don't know how to set it up in system of equations

OpenStudy (anonymous):

but i can solve it perfectly fine

OpenStudy (anonymous):

it says to use two equation (which is not really necessary) but we can do it if \(x\) is the amount of the investment that lost 3% and y is the investment that gained 7% then one equation is \[x+y=11,500\]

OpenStudy (anonymous):

the other would be \[-.03x+.07y=196\]

OpenStudy (anonymous):

solve the "system" \[x+y=11,500\\ -.03x+.07y=196\] which might be easier if you multiplied the second one by 100 and had \[x+y=11500\\ -3x+7t=19,600\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

@satellite73 yes thank you so much!!!

OpenStudy (anonymous):

yw